A drug store buys 2,000 rolls of Christmas paper for $3.00 per roll. It is expected that 10% of the paper will eventually be sold for $2.25 per roll. If the store wishes to make a gross profit of 50% on cost, what should be the marked price of a roll of paper?

a. $4.50
b. $6.67
c. $4.75
d. $3.75


Answer by ankor@dixie-net.com(22740)

A drug store buys 2,000 rolls of Christmas paper for $3.00 per roll. It is expected that 10% of the paper will eventually be sold for $2.25 per roll. If the store wishes to make a gross profit of 50% on cost, what should be the marked price of a roll of paper?
:
Total cost of the paper: 2000 * 3 = $6000
:
.10 * 2000 = 200 rolls will be sold at $2.25,
Revenue from these: 200 * 2.25 = $450
:
To make 50% on the cost: .5 * $6000 = $3000
:
Let x = price required to make 50% profit
:
x(2000 - 200) + 450 - 6000 = 3000
1800x - 5550 = 3000
1800x = 3000 + 5550
1800x = 8550
x = 8550/1800
x = $4.75